Calculate the mass moment of inertia for the components and axes given below: a) A 3 kg rod which is 1.5m long about its centre b) A 2kg metal plate 300mmx400mm about an axis perpendicular to the plane of the plate through its centre. c) A 500g disc of diameter 100mm about an axis perpendicular to the plane of the plate through its centre. 


1
Expert's answer
2021-07-20T07:09:02-0400

a.

"I= \frac{m}{12} * L\\
I= \frac{3}{12} * 1.5\\
I= 0.5625 kgm^2\\"

b.

"I_x= \frac{M}{12} *a^2\\
I_x= \frac{2}{12} *(300*300)\\
I_x= 15000000 gmm^2\\
I_y= \frac{M}{12} *b^2\\
I_y= \frac{2}{12} *(400*400)\\
I_y= 26666666.666667 gmm^2\\
I_z= \frac{1}{12}*M*(a^2+b^2)\\
I_z= \frac{1}{12}*2*(300^2+400^2)
I_z=41666666.666667 gmm^2"

c.

"I_{y} = m * r²\\
I_{y} = 500 * 50²\\
I_{y} =0.0003125 kgm^2\\

I_{x} = m * r²\\
I_{x} = 500 * 50²\\
I_{x} =0.0003125 kgm^2\\

I_{z} = 0.5m * r²\\
I_{z} = 0.5*500 * 50²\\
I_{z} =0.000625 kgm^2\\"